OpenStudy (anonymous):

derivative of 5x^cos (x^x)

OpenStudy (anonymous):

****5x^cos(x^2)

OpenStudy (anonymous):

Take the logarithm of both sides first, and then apply the derivative operator:\[y=5x^{\cos x^2} \rightarrow \ln y=\ln 5+\cos (x^2) \ln x\]

OpenStudy (anonymous):

Then\[\frac{y'}{y}=\frac{\cos(x^2)}{x}+\ln x \times -\sin(x^2) \times 2x\]

OpenStudy (anonymous):

Multiply both sides by y = 5x^(cos(x^2)) to get\[y'=5x^{\cos{x^2}}\left[ \frac{\cos (x^2)}{x}-2x \sin (x^2) \ln x \right]\]

OpenStudy (anonymous):

Just check my working...rushing...